If a=(xz^(3))i-(2x^(2)yz)j+(2yz^(4))k find curl a at (1 -1 1)


1
Expert's answer
2021-01-26T04:38:31-0500
"\vec a=xz^3\vec i-2x^2yz\vec j+2yz^4\vec k"

"\text{curl } \vec a=\nabla\times\vec a=\begin{vmatrix}
 \vec i & \vec j & \vec k \\ \\
 \dfrac{\partial}{\partial x} & \dfrac{\partial}{\partial y} & \dfrac{\partial}{\partial z} \\ \\
xz^3& -2x^2yz & 2yz^4
\end{vmatrix}"

"=\vec i(2z^4-(-2x^2y)-\vec j(0-3xz^2)+\vec k(-4xyz-0)"

"=(2z^4+2x^2y)\vec i+3xz^2\vec j-4xyz\vec k"

"\text{curl } \vec a\big|_{(1,-1,1)}=3\vec j+4\vec k"